OpenStudy (kaylaprincess):

I don't understand how to get the answer because the equation is funky? What are the x-coordinates of the solutions to this system of equations? x^2 + y^2 = 25 y = 2x - 5

OpenStudy (kaylaprincess):

I know to plug in y into x^2 + y^2 = 25, im just confused because the square i think?

OpenStudy (phi):

the first equation is a circle the second is a line they are asking where the line crosses the circle.

OpenStudy (kaylaprincess):

so all i needed to do was graph it lol

OpenStudy (phi):

yes, it is a bit ugly but y= 2x-5 so you can replace y in the first equation \[ x^2 + (2x-5)^2 = 25\]

OpenStudy (kaylaprincess):

this is a bit further question, but could i just square root the equation and get rid of the squares?

OpenStudy (phi):

no, the square root of a sum is ugly you have to expand (2x-5)(2x-5) as the first step

OpenStudy (kaylaprincess):

then plug in y?

OpenStudy (kaylaprincess):

ohh. okay. well thank you as always.

OpenStudy (phi):

you should get (2x-5)(2x-5)= 4x^2 -20x + 25 so the equation is now \[ x^2 + 4x^2 -20x + 25 = 25\]

OpenStudy (phi):

and here is what the problem looks like when graphed